# Popov Mirror-Prox for solving Variational Inequalities

### v1

@@ I Introduction
Variational Inequalities (VIs) are important for the study of equilibrium problems arising in economics [1], game theory [2], multi-agent reinforcement learning [3], etc. To solve the VI problems, conventional projection-based methods have been developed in [4, 5, 6, 7, 8]. Korpelevich [5] and Popov [6] methods have been mostly used to solve VI problems with monotone mappings, as the projection method can diverge for such operators [9].
The mirror descent algorithm was introduced in [10] in the context of optimization problems, and has dimension (of the variable) independent rates for non-smooth functions [11, 12]. It has been also shown to have a favorable scaling with the gradient size [13] in comparison with other projection-based first-order methods. The mirror-prox algorithm in the context of monotone VIs has been proposed in [12] for the deterministic VIs and in [14] for the stochastic VIs. However, a direct implementation of the mirror-prox method is rarely possible. To address this issue, the work in [12] has proposed the use of Korpelevich method (aka extra-gradient method) to develop an implementable version of the mirror-prox method. The Korpelevich mirror-prox algorithm requires two mapping evaluations per iteration. Instead, a single mapping evaluation per iteration can be used if the Popov method is employed in mirror-prox as proposed in [15] for pseudo-monotone VIs. However, the convergence rate of the algorithm has remained an open question. On the other hand, the work in [16] presents the optimistic/Popov methods for saddle point problems and provide convergence rates. The work in [17] studies similar algorithms but with strongly monotone mapping. All of the aforementioned works, do not provide convergence rate of the Popov mirror-prox method for monotone VIs, which we address in this paper.
@@ II-A The VI problem
When the set $X$ is convex and compact, and the mapping $F(\cdot)$ is monotone on $X$ (cf. Assumptions 12) a weak solution always exists [24]. Moreover, when the mapping $F(\cdot)$ is monotone and continuous on $X$, a weak4 solution is also a strong solution [24]. Under Assumption 3, the mapping $F(\cdot)$ is continuous. Hence, under Assumptions 13, the weak and the strong solutions of the VI$(X,F)$ exist and coincide.
@@ II-B Preliminaries and Mirror-prox Algorithm
These functions will be important in the subsequent analysis of the Popov mirror-prox method. Moreover, we will use the prox-mapping defined as follows: For every $x\in X$ and $\xi\in E^{*}$, the prox-mapping $P_{x}:E^{*}\rightarrow X$ is given by
@@
Thus, determining the point $x_{t}$ in a closed form is not always easy, unless the set $X$ has a favorable structure so that it is easy to project on. In this paper, we will assume that the set $X$ has a simple structure for the projection, and we will be focusing only on simplifying the update for $y_{t}$ so that inequality (10) is satisfied.
@@
where $L$ is the Lipschitz constant of the mapping $F(\cdot)$ from Assumption 3. With a choice of the step size $\gamma$, such that $\gamma L/\alpha<1$, from (12) we see that the prox-mapping is contractive (at points of the form $\gamma F(y)$) and, thus, the solution $x_{t}$ in (11) can be attained geometrically fast. It turned out that $x_{t}$ can be obtained with only two update steps (see [12]), and we consider such updates in the next section III for the mirror-prox method combined with the Popov method.
@@ III Popov Mirror-prox Method
In Algorithm 2, the updates for $y_{t+1}$ and $x_{t+1}$ are obtained via relations in (13), by letting $x=x_{t}$, $\widetilde{y}=y_{t+1}$, $\widetilde{x}=x_{t+1}$, $\xi=F(y_{t})$, $\eta=F(y_{t+1})$, and using a fixed step size $\gamma>0$. Since both $y_{t+1}$ and $x_{t+1}$ at step 3 of Algorithm 2 are solutions to strongly convex problems, they are uniquely defined. Our assumption is that the set $X$ and the function $\psi(\cdot)$ has a simple structure so that the iterates $y_{t+1}$ and $x_{t+1}$ have a closed form expression.
Moreover, by using the optimality conditions, we can see that the iterates $y_{t+1}$ and $x_{t+1}$ are such that for any $z\in X$, the following inequalities hold:
@@
The first step of the update produces an iterate sequence of $y_{t+1}$’s and the next step produces another iterate sequence of $x_{t+1}$’s. The mapping $F(\cdot)$ is evaluated with respect to the sequence $\{y_{t}\}$ and the old mapping is reused resulting in a single mapping computation per loop of the algorithm. The output $y^{(t)}$ of Algorithm 2 is the average of the iterates $y_{k}$, $12\leq k\leq t+$ 1.
@@ IV Convergence Rate analysis
Lemma 2 will be used to establish the convergence rate of Algorithm 2. Note that the first two equations of Lemma 2 has a lot of resemblance with (12) and the relations presented in Lemma 1. The relation (17) is useful for analyzing the convergence rates of Algorithm 2.
@@
Next, we present a lemma that establishes an upper bound for the sum of $\epsilon_{t}$ for index $t$ ranging from $1$ to some $T\geq 1$.
@@
Let Assumptions 1, 3, and 4 hold. Consider Algorithm 2 with the constant step size $0<\gamma\leq\frac{\alpha}{2L}$. Then, we have for all $T\geq 1$,
$\sum_{t=1}^{T}\epsilon_{t}\leq$ $\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{1}-x_{1}\|^{2}.$
@@
Next, we sum of all $\epsilon_{t}$ for $t=1,\ldots,T$, and obtain
$\sum_{t=1}^{T}\epsilon_{t}$ $\leq\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{1}-x_{1}\|^{2}-\frac{\alpha}{2}\|y_{T+1}-x_{T+1}\|^{2}$
$-\sum_{t=1}^{T-1}\left(\frac{\alpha}{2}-\frac{2\gamma^{2}L^{2}}{\alpha}\right)\|y_{t+1}-x_{t+1}\|^{2}$
$-\sum_{t=0}^{T-1}\left(\frac{\alpha}{2}-\frac{2\gamma^{2}L^{2}}{\alpha}\right)\|y_{t+1}-x_{t}\|^{2}.$ (22)
@@
Let Assumptions 14 hold. Consider Algorithm 2 with a constant step size $\gamma\leq\frac{\alpha}{2L}$. Then, for the dual gap function evaluated at the average $y^{(}T)$ of the iterates iterates, we have for all $T\geq 1$,
$\max_{u\in X}\langle F(u),y^{(T)}\!-\!u\rangle\leq\frac{2L}{T\alpha}\max_{u\in X}B_{\psi}(u,x_{1})\!+\!\frac{L}{T}\|y_{1}-x_{1}\|^{2},$
where $y^{(T)}=\frac{1}{T}\sum_{t=1}^{t}y_{t+1}$ and $B_{\psi}(\cdot,\cdot)$ is the Bregman divergence induced by $\psi(\cdot)$, defined by $B_{\psi}(u,x)=\psi(u)-\psi(x)-\langle\nabla\psi(x),u-x\rangle$ for all $u,x\in X$.
@@
We start the proof by using (17) of Lemma 2 with the follwoing identifications $x=x_{t}$, $\widetilde{y}=y_{t+1}$, $\widetilde{x}=x_{t+1}$, $\xi=F(y_{t})$, $\eta=F(y_{t+1})$, $\delta=\delta_{t}$, and $\epsilon=\epsilon_{t}$. We thus obtain
@@
By summing the preceding relation for $t=1,\ldots,T$, for any $T\geq 0$, we obtain
$\sum_{t=1}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle\leq H_{u}(x_{1})-H_{u}(x_{T+1})+\sum_{t=1}^{T}\epsilon_{t}.$ (24)
@@
$\sum_{t=1}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle$ $\geq\sum_{t=1}^{T}\langle\gamma F(u),y_{t+1}-u\rangle$
$=\gamma\sum_{t=1}^{T}\langle F(u),y_{t+1}-u\rangle$
$=\langle F(u),\sum_{t=1}^{T}y_{t+1}-u\rangle.$
Since $\sum_{t=1}^{T}y_{t+1}=Ty^{(T)}$, it follows that
$\sum_{t=1}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle\geq\gamma T\langle F(u),y^{(T)}-u\rangle.$ (25)
@@
$H_{u}(x_{1})$ $=\Omega(\nabla\psi(x_{1}))-\langle\nabla\psi(x_{1}),u\rangle$
$=\max_{x\in X}[\langle\nabla\psi(x_{1}),x\rangle-\psi(x)]-\langle\nabla\psi(x_{1}),u\rangle.$ (26)
@@
$\langle\nabla\psi(x_{1})-\nabla\psi(\hat{x}),x-\hat{x}\rangle\leq 0\qquad\forall x\in X.$
Note that $x_{1}\in X$, i.e., it is feasible based on the updates of Algorithm 2 and, hence, we see that the solution is $\hat{x}=x_{1}$, as the preceding inequality is satisfied. Therefore, by using this in the right hand side of (26), we obtain
$H_{u}(x_{1})=\langle\nabla\psi(x_{1}),x_{1}-u\rangle-\psi(x_{1}).$ (27)
In the same way, we can derive the relation for the second term on the right hand side of (24) as
@@
$\gamma T\langle F(u),y^{(T)}-u\rangle$ $\leq\psi(u)-\psi(x_{1})-\langle\nabla\psi(x_{1}),u-x_{1}\rangle$
$-\frac{\alpha}{2}\|u-x_{T+1}\|^{2}+\sum_{t=1}^{T}\epsilon_{t}.$ (29)
Now the quantity $\psi(u)-\psi(x_{1})-\langle\nabla\psi(x_{1}),u-x_{1}\rangle$ on the right hand side of (29) is denoted by $B_{\psi}(u,x_{1})$ which represents the Bregman divergence between the point $u\in X$ and $x_{1}\in X$ induced by the function $\psi$. In addition, the last term on the right hand side of (29) can be upper bounded using Lemma 3. With these steps, (29) reduces to
$\gamma T\langle F(u),y^{(T)}-u\rangle+\frac{\alpha}{2}\|u-x_{T+1}\|^{2}$
$\leq B_{\psi}(u,x_{1})+\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{1}-x_{1}\|^{2}.$
In the preceding relation, we can ignore the second term on the left hand side and take the maximum with respect to $u\in X$ on both sides of the relation. Thus, we obtain
$\gamma T\max_{u\in X}\langle F(u),y^{(T)}-u\rangle\leq\max_{u\in X}B_{\psi}(u,x_{1})+\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{1}-x_{1}\|^{2}.$
Diving by $\gamma T$ and using the dual gap function (see (5), we have that
$G(y^{(T)})\leq\frac{1}{T\gamma}\max_{u\in X}B_{\psi}(u,x_{1})+\frac{2\gamma L^{2}}{\alpha T}\|y_{1}-x_{1}\|^{2}.$
The result follows by using $\gamma\leq\frac{\alpha}{2L}$.  Q.E.D.
Theorem 1 shows $O(1/T)$ rate of convergence of the dual gap function for the Popov mirror-prox which is the same rate as that of Korpelevich mirror-prox studied in [12], but with an additional term of $\frac{L}{T}\|y_{1}-x_{1}\|^{2}$ coming from the upper bound on $\sum_{t=1}^{T}\epsilon_{t}$ in Lemma 3. Moreover, since the set $X$ is compact by Assumption 1, the quantity $\max_{u\in X}B_{\psi}(u,x_{1})$ is finite.
@@ V Simulations
In this section we present simulation results for Popov and Korpelevich versions of the mirror-prox algorithm on a two player matrix game given by
@@
where $\Delta_{2}$ is the probability simplex in $\mathbb{R}^{2}$, i.e., $\Delta_{2}=\{(h_{1},h_{2})\mid h\succcurlyeq 0,\ h_{1}+h_{2}=1\}$. Thus, Assumption 1 is satisfied. The game can be modeled as a VI problem with the mapping $F(\cdot)$ obtained by differentiating each agent’s loss function with respect to their respective decision variables, i.e.,
@@
where $x=[x_{1},x_{2}]^{T}\in\mathbb{R}^{4}$. Thus, the two player matrix game problem is equivalent to VI$(X,F)$ with $X=\Delta_{2}\times\Delta_{2}$ and the mapping $F(\cdot)$ as in (31). The matrices $A$, $B$, $C$, and $D$ are generated to make the Jacobian of $F(\cdot)$, denoted by $\nabla F$, positive semidefinite with eigenvalues in the interval $[0,100]$. Hence, both Assumptions 2 and 3 are satisfied with $L=100$. The vectors $p$ and $s$ are sampled from a random normal distribution. For more details about the generation process see [26, Appendix F.1]. We experiment with two different functions $\psi(\cdot)$: (i) Entropic case: $\psi(x)=\sum_{j=1}^{4}x_{j}(\ln x_{j}-1)$ implies $\nabla\psi(x)=[\ln x_{1},\ldots,\ln x_{4}]^{T}$. (ii) Euclidean case: $\psi(x)=\frac{1}{2}\|x\|^{2}$ implies $\nabla\psi(x)=x$. For the entropic case, $\nabla^{2}\psi(x)\succcurlyeq\frac{I}{\max_{i}x_{i}}\succcurlyeq I$, since $\max_{i}x_{i}\leq 1$ due to the probability simplex constraints, where $I$ is the identity matrix. For the Euclidean case, $\nabla^{2}\psi(x)=I$. Thus, in both cases, Assumption 4 is satisfied with $\alpha=1$. We run Algorithm 2 using the step size $\gamma=\frac{\alpha}{2L}=\frac{1}{2L}$ and compare it with the Korpelevich mirror-prox algorithm [12] with the prescribed step size $\gamma=\frac{\alpha}{\sqrt{2}L}$. For the entropic case, the inverse transformation $\nabla\psi(\xi)^{-1}=\exp(\xi)$ preserves the nonnegativity of the variables. Therefore, the iterates $y_{t+1}$ and $x_{t+1}$ at step 3 of Algorithm 2 are given by
@@
The dual gap function cannot be evaluated exactly and we estimate it empirically by generating $50000$ random samples within the simplex $\Delta_{2}\times\Delta_{2}$ and determining the maximum value of the inner product in (5). Note that empirical approximation can be erroneous and introduce a variance between empirical and actual dual gap function depending on the number of the samples generated. To lower the estimation variance, we considered the matrix game on dimension 2.
@@
Fig. 1 shows two different experiments with randomly generated Jacobian $\nabla F(\cdot)$. It can be seen that for the entropic $\psi$ the results are better than for the Euclidean $\psi$ in the first plot of Fig. 1 for both Korpelevich and Popov mirror-prox methods. In the second plot, the results for the Euclidean $\psi$ are better than for the entropic $\psi$ for both methods. In both plots, the Korpelevich mirror-prox has better performance than that of the Popov mirror-prox. We believe this is due to a larger step size used for the former and an additional error term $\frac{L}{T}\|y_{1}-x_{1}\|^{2}$ in the analysis of the later. But, the Popov mirror-prox requires a single mapping computation, and hence it is computationally less expensive compared to the Korpelevich mirror-prox.

### v2

@@ I Introduction
Variational Inequalities (VIs) are important for the study of equilibrium problems arising in economics [1], game theory [2], multi-agent reinforcement learning [3], etc. To solve VI problems, conventional projection-based methods have been developed in [4, 5, 6, 7, 8]. Korpelevich [5] and Popov [6] methods have been mostly used to solve VI problems with monotone mappings, as the projection method can diverge for such mappings [9].
The mirror descent algorithm was introduced in [10] in the context of optimization problems, and has dimension (of the variable) independent rates for non-smooth functions [11, 12]. It has been also shown to have a favorable scaling with the gradient size [13] in comparison with other projection-based first-order methods. The mirror-prox algorithm in the context of monotone VIs has been proposed in [12] for the deterministic VIs and in [14] for the stochastic VIs. However, a direct implementation of the mirror-prox method is rarely possible. To address this issue, the work in [12] has proposed the use of Korpelevich method (aka extra-gradient method) to develop an implementable version of the mirror-prox method. The Korpelevich mirror-prox algorithm requires two mapping evaluations per iteration. Instead, a single mapping evaluation per iteration can be used if the Popov method is employed in mirror-prox as proposed in [15] for pseudo-monotone VIs. However, the convergence rate of the algorithm has remained an open question. On the other hand, the work in [16] presents the optimistic/Popov methods for saddle point problems and provide convergence rates. The work in [17] studies similar algorithms but with strongly monotone mapping. None of the aforementioned work provides convergence rate of the Popov mirror-prox method for monotone VIs, which we address in this paper.
@@ II-A VI Problem and Assumptions
When the set $X$ is convex and compact, and the mapping $F(\cdot)$ is monotone on $X$ (cf. Assumptions 12) a weak solution always exists [24]. Moreover, when the mapping $F(\cdot)$ is monotone and continuous on $X$, a weak solution is also a strong solution [24]. Under Assumption 3, the mapping $F(\cdot)$ is continuous. Hence, under Assumptions 13, the weak and the strong solutions of the VI$(X,F)$ exist and coincide.
@@ II-B Preliminaries and Mirror-prox Algorithm
These functions will be important in the subsequent analysis of the Popov mirror-prox method. The function $H_{u}(x)$ captures a form of distance between the points $u\in X$ and $x\in X$, and it is an artifact of the analysis. Moreover, we will use the prox-mapping defined as follows: For every $x\in X$ and $\xi\in E^{*}$, the prox-mapping $P_{x}:E^{*}\rightarrow X$ is given by
@@
Thus, determining the point $x_{t}$ in a closed form is not always easy, unless the set $X$ has a favorable structure so that it is easy to project on, such as, for example box constraints, norm ball, etc. In this paper, we assume that the set $X$ has a simple structure for the projection, and we focus only on simplifying the update for $y_{t}$ so that inequality (10) is satisfied.
@@
where $L$ is the Lipschitz constant of the mapping $F(\cdot)$ from Assumption 3. With a choice of the step size $\gamma$, such that $\gamma L/\alpha<1$, from (12) we see that the prox-mapping is contractive (at points of the form $\gamma F(y)$) and, thus, the solution $x_{t}$ in (11) can be obtained geometrically fast. It turned out that $x_{t}$ can be obtained with only two update steps [12]. We consider such updates in the next section for the mirror-prox method combined with the Popov method.
@@ III Popov Mirror-prox Method
In Algorithm 2, the updates for $y_{t+1}$ and $x_{t+1}$ are obtained via relations in (13), by letting $x=x_{t}$, $\widetilde{y}=y_{t+1}$, $\widetilde{x}=x_{t+1}$, $\xi=F(y_{t})$, $\eta=F(y_{t+1})$, and using a fixed step size $\gamma>0$. Since both $y_{t+1}$ and $x_{t+1}$ at step 3 of Algorithm 2 are solutions to strongly convex problems, they are uniquely defined. Moreover, by using the optimality conditions, we can see that the iterates $y_{t+1}$ and $x_{t+1}$ are such that for any $z\in X$, the following inequalities hold:
@@
The first step of the update produces an iterate sequence $\{y_{t}\}$, while the next step produces another iterate sequence $\{x_{t}\}$. The mapping $F(\cdot)$ is evaluated with respect to the sequence $\{y_{t}\}$ and the old mapping is reused resulting in a single mapping computation per loop of the algorithm. The output $y^{(t+1)}$ of Algorithm 2 is the average of the iterates $y_{k}$, $11\leq k\leq t+$ 1.
@@ IV Convergence Rate analysis
Lemma 2 will be used to establish the convergence rate of Algorithm 2. Note that the first two equations of Lemma 2 have a lot of resemblance with (12) and the relations presented in Lemma 1. The relation in (17) is useful for analyzing the convergence rate of Algorithm 2.
@@
Next, we present a lemma that establishes an upper bound for the sum of $\epsilon_{t}$ for $t$ ranging from $1$ to some $T\geq 1$.
@@
Let Assumptions 1, 3, and 4 hold. Consider Algorithm 2 with a constant step size $\gamma$ satisfying $0<\gamma\leq\frac{\alpha}{2L}$. Then, we have for all $T\geq 1$,
$\sum_{t=0}^{T}\epsilon_{t}\leq\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{0}-x_{0}\|^{2}.$
@@
Next, we sum $\epsilon_{t}$ for $t=0,1,\ldots,T$, and obtain
$\sum_{t=0}^{T}\epsilon_{t}$ $\leq\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{0}-x_{0}\|^{2}-\frac{\alpha}{2}\|y_{T+1}-x_{T+1}\|^{2}$
$-\sum_{t=0}^{T-1}\left(\frac{\alpha}{2}-\frac{2\gamma^{2}L^{2}}{\alpha}\right)\|y_{t+1}-x_{t+1}\|^{2}$
$-\sum_{t=0}^{T}\left(\frac{\alpha}{2}-\frac{2\gamma^{2}L^{2}}{\alpha}\right)\|y_{t+1}-x_{t}\|^{2}.$ (22)
@@
Let Assumptions 14 hold. Consider Algorithm 2 with a constant step size $\gamma=\frac{\alpha}{2L}$. Then, for the dual gap function $G(\cdot)$, evaluated at the average $y^{(T+1)}$ of the iterates, we have for all $T\geq 1$,
$G(y^{(T+1)})\leq\frac{2L}{(T+1)\alpha}\max_{u\in X}B_{\psi}(u,x_{0})+\frac{L}{T+1}\|y_{0}-x_{0}\|^{2},$
where $y^{(T+1)}=\frac{1}{T+1}\sum_{t=0}^{T}y_{t+1}$ and $B_{\psi}(\cdot,\cdot)$ is the Bregman divergence induced by $\psi(\cdot)$, defined by $B_{\psi}(u,x)=\psi(u)-\psi(x)-\langle\nabla\psi(x),u-x\rangle$ for all $u,x\in X$.
@@
We use (17) of Lemma 2 with the following identifications $x=x_{t}$, $\widetilde{y}=y_{t+1}$, $\widetilde{x}=x_{t+1}$, $\xi=F(y_{t})$, $\eta=F(y_{t+1})$, $\delta=\delta_{t}$, and $\epsilon=\epsilon_{t}$. We thus obtain
@@
By summing the preceding relation for $t=0,1,\ldots,T$, for any $T\geq 1$, we obtain
$\sum_{t=0}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle\leq H_{u}(x_{0})-H_{u}(x_{T+1})+\sum_{t=0}^{T}\epsilon_{t}.$ (24)
@@
$\sum_{t=0}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle\geq\sum_{t=0}^{T}\langle\gamma F(u),y_{t+1}-u\rangle$
Since $\sum_{t=0}^{T}y_{t+1}=(T+1)y^{(T+1)}$, it follows that
$\sum_{t=0}^{T}\langle\gamma F(y_{t+1}),y_{t+1}-u\rangle\geq\gamma(T+1)\langle F(u),y^{(T+1)}-u\rangle.$ (25)
@@
$H_{u}(x_{0})$ $=\Omega(\nabla\psi(x_{0}))-\langle\nabla\psi(x_{0}),u\rangle$
$=\max_{x\in X}[\langle\nabla\psi(x_{0}),x\rangle-\psi(x)]-\langle\nabla\psi(x_{0}),u\rangle.$ (26)
@@
$\langle\nabla\psi(x_{0})-\nabla\psi(\hat{x}),x-\hat{x}\rangle\leq 0\qquad\forall x\in X.$
Note that $x_{0}\in X$, i.e., it is feasible based on the updates of Algorithm 2 and, hence, we see that the solution is $\hat{x}=x_{0}$, as the preceding inequality is satisfied. Therefore, by using this in the right hand side of (26), we obtain
$H_{u}(x_{0})=\langle\nabla\psi(x_{0}),x_{0}-u\rangle-\psi(x_{0}).$ (27)
In the same way, we can derive the relation for the second term on the right hand side of (24), and obtain
@@
$\gamma(T+1)\langle F(u),y^{(T+1)}-u\rangle\leq\psi(u)-\psi(x_{0})$
$-\langle\nabla\psi(x_{0}),u-x_{0}\rangle-\frac{\alpha}{2}\|u-x_{T+1}\|^{2}+\sum_{t=0}^{T}\epsilon_{t}.$ (29)
Now the quantity $\psi(u)-\psi(x_{0})-\langle\nabla\psi(x_{0}),u-x_{0}\rangle$ on the right hand side of (29) is denoted by $B_{\psi}(u,x_{0})$ which represents the Bregman divergence between the point $u\in X$ and $x_{0}\in X$ induced by the function $\psi$. In addition, the last term on the right hand side of (29) can be upper bounded using Lemma 3. With these steps, (29) reduces to
$\gamma(T+1)\langle F(u),y^{(T+1)}-u\rangle+\frac{\alpha}{2}\|u-x_{T+1}\|^{2}$
$\leq B_{\psi}(u,x_{0})+\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{0}-x_{0}\|^{2}.$
In the preceding relation, we can ignore the second term on the left hand side and take the maximum with respect to $u\in X$ on both sides of the relation. Additionally, using the dual gap function (see (5)), we obtain
$\gamma(T+1)G(y^{(T+1)})\leq\max_{u\in X}B_{\psi}(u,x_{0})+\frac{2\gamma^{2}L^{2}}{\alpha}\|y_{0}-x_{0}\|^{2}.$
Dividing by $\gamma(T+1)$ both sides of the preceding relation, we obtain
$G(y^{(T+1)})\leq\frac{\max_{u\in X}B_{\psi}(u,x_{0})}{(T+1)\gamma}+\frac{2\gamma L^{2}}{(T+1)\alpha}\|y_{0}-x_{0}\|^{2}.$
The result follows by using $\gamma=\frac{\alpha}{2L}$. Moreover, since the set $X$ is compact by Assumption 1, the quantity $\max_{u\in X}B_{\psi}(u,x_{0})$ is finite.  Q.E.D.
Theorem 1 shows $O(1/T)$ convergence rate of the dual gap function for the Popov mirror-prox which is the same as that of Korpelevich mirror-prox studied in [12], but with an additional term of $\frac{L}{T+1}\|y_{0}-x_{0}\|^{2}$ coming from the upper bound on $\sum_{t=0}^{T}\epsilon_{t}$ in Lemma 3. We can eliminate the additional term by initializing Algorithm 2 with $x_{0}=y_{0}\in X$, hence giving a rate result similar to that of the Korpelevich method but with only a half of the mapping evaluations.
@@ V Simulations
In this section, we present simulation results for Popov and Korpelevich variants of the mirror-prox algorithm on a two player matrix game given by
@@
where $\Delta_{2}$ is the probability simplex in $\mathbb{R}^{2}$, i.e., $\Delta_{2}=\{(h_{1},h_{2})\mid h\succcurlyeq 0,\ h_{1}+h_{2}=1\}$. Hence, Assumption 1 is satisfied. The game can be modeled as a VI problem with the mapping $F(\cdot)$ obtained by differentiating each agent’s loss function with respect to their respective decision variables, i.e.,
@@
where $x=[x_{1},x_{2}]^{T}\in\mathbb{R}^{4}$. Thus, the two player matrix game problem is equivalent to VI$(X,F)$ with $X=\Delta_{2}\times\Delta_{2}$ and the mapping $F(\cdot)$ as in (31). The matrices $A$, $B$, $C$, and $D$ are generated to make the Jacobian of $F(\cdot)$, denoted by $\nabla F$, positive semidefinite with eigenvalues in the interval $[0,100]$. Hence, both Assumptions 2 and 3 are satisfied with $L=100$. The vectors $p$ and $s$ are sampled from a random normal distribution. For more details about the generation process see [26, Section 7]. We experiment with two different functions $\psi(\cdot)$: (i) Entropic case: $\psi(x)=\sum_{j=1}^{2}\sum_{i=1}^{2}x_{j}^{(i)}(\ln x_{j}^{(i)}-1)$, which implies $\nabla\psi(x)=[\ln x_{1}^{(1)},\ln x_{1}^{(2)},\ln x_{2}^{(1)},\ln x_{2}^{(2)}]^{T}$, where $x_{j}=[x_{j}^{(1)},x_{j}^{(2)}]\in\mathbb{R}^{2}$ for $j=1,2$. (ii) Euclidean case: $\psi(x)=\frac{1}{2}\|x\|^{2}$, which implies $\nabla\psi(x)=x$. For the entropic case, $\nabla^{2}\psi(x)\succcurlyeq\frac{I}{\max_{j,i}x_{j}^{(i)}}\succcurlyeq I$, since $\max_{j,i}x_{j}^{(i)}\leq 1$ due to the probability simplex constraints, where $I$ is the identity matrix. For the Euclidean case, $\nabla^{2}\psi(x)=I$. Thus, in both cases, Assumption 4 is satisfied with $\alpha=1$. We run Algorithm 2 using the step size $\gamma=\frac{\alpha}{2L}=\frac{1}{2L}$ and compare it with the Korpelevich mirror-prox algorithm [12] with the prescribed step size $\gamma=\frac{\alpha}{\sqrt{2}L}$. For the entropic case, the inverse transformation $\nabla\psi(\xi)^{-1}=\exp(\xi)$ preserves the nonnegativity of the variables. Therefore, the iterates $y_{t+1}$ and $x_{t+1}$ at step 3 of Algorithm 2 are given by
@@
We initialize Algorithm 2 with $x_{0}=y_{0}=[0.5,0.5]^{2}$. The dual gap function cannot be evaluated exactly and we estimate it empirically by generating $200000$ random samples within the simplex $\Delta_{2}\times\Delta_{2}$ and determining the maximum value of the inner product in (5). Note that this empirical approximation can be erroneous and introduce a variance between the empirical and the actual dual gap functions depending on the number of the samples generated. To lower the variance, we considered the matrix game of dimension 2.
Fig. 1 shows two different experiments with randomly generated Jacobian $\nabla F(\cdot)$. The left plot of Fig. 1 shows a superior performance for the entropic $\psi$ to that of the Euclidean $\psi$, whereas in the right plot of Fig. 1, it is vice versa for both Korpelevich and Popov mirror-prox methods. Additionally, in both plots, the performance of Korpelevich and Popov mirror-prox methods are similar since we have chosen the same $x_{0}$ and $y_{0}$ for the later in order to eliminate the error term $\frac{L}{T}\|y_{0}-x_{0}\|^{2}$ showing up in the analysis. Moreover, Popov mirror-prox is computationally less expensive as compared to the Korpelevich mirror-prox.